21. Graph the following polar curves: a) r = 2 + 4 cos(O) b) r = 5 – 10 sin(0) For each of the graphs above, set up the integonly do part (a)
Answer #1

(a) r = 2+4cos(O) r=0=2+4 cos() = 0 cos(2) 2 27 47 3 3 area of inner loop: $ $(2+4cos (0))*d@= 481–623 area of outer loop: us